### v1

# On full-separating sets in graphs

Abstract
Several different types of identification problems have been already studied in the literature, where the objective is to distinguish any two vertices of a graph by their unique neighborhoods in a suitably chosen dominating or total-dominating set of the graph, often referred to as a code. To study such problems under a unifying point of view, reformulations of the already studied problems in terms of covering problems in suitably constructed hypergraphs have been provided. Analyzing these hypergraph representations, we introduce a new separation property, called full-separation, which has not yet been considered in the literature so far. We study it in combination with both domination and total-domination, and call the resulting codes full-separating-dominating codes (or FD-codes for short) and full-separating-total-dominating codes (or FTD-codes for short), respectively. We address the conditions for the existence of FD- and FTD-codes, bounds for their size and their relation to codes of the other types. We show that the problems of determining an FD- or an FTD-code of minimum cardinality in a graph is NP-hard. We also show that the cardinalities of minimum FD- and FTD-codes differ by at most one, but that it is NP-complete to decide if they are equal for a given graph in general. We find the exact values of minimum cardinalities of the FD- and FTD-codes on some familiar graph classes like paths, cycles, half-graphs and spiders. This helps us compare the two codes with other codes on these graph families thereby exhibiting extremal cases for several lower bounds.
Keywords:
full-separation open-separation closed-separation domination total-domination NP-completeness hypergraph
1 Introduction
In the domain of identification problems, the objective is typically to distinguish any two vertices of a graph by the unique intersection of their neighborhoods with a suitably chosen dominating or total-dominating set of the graph. Depending on the types of dominating sets and separation properties used, various problems arise under various names in the literature.
More precisely, consider a graph $G=(V,E)$ and denote by $N(v)=\{u\in V:uv\in E\}$ (respectively, $N[v]=N(v)\cup\{v\}$) the open (respectively, closed) neighborhood of a vertex $v\in V$. On the one hand, a subset $C\subseteq V$ is dominating (respectively, total-dominating) if the set $N[v]\cap C$ (respectively, $N(v)\cap C$) is non-empty for each $v\in V$. On the other hand, a subset $C\subseteq V$ is closed-separating (respectively, open-separating) if the set $N[v]\cap C$ (respectively, $N(v)\cap C$) is unique for each $v\in V$. Moreover, $C$ is locating if the set $N(v)\cap C$ is unique for each $v\in V\setminus C$. So far, the following combinations of location/separation and domination properties have been studied in the literature:
  • closed-separation with domination and total-domination leading to identifying codes (ID-codes for short) and identifying-total-dominating codes [^1] Identifying-total-dominating codes had been introduced to the literature in [13] under the name differentiating total-dominating codes. However, due to consistency in notation, we prefer to call them ITD-codes in this article. (ITD-codes for short), see [16] and [13], respectively;
  • location with domination and total-domination leading to locating-dominating codes (LD-codes for short) and locating-total-dominating codes (LTD-codes for short), see [23] and [13], respectively;
  • open-separation with domination and total-domination leading to open-separating-dominating codes (OD-codes for short) and open-separating-total-dominating codes [^2] OTD-codes were first introduced as open neighborhood locating-dominating sets and are also popularly known as OLD-sets. However, due to consistency in notation, we prefer to call them OTD-codes in this article. Moreover, OTD-codes were also introduced in a more general setting in [14] under the concepts of strongly $(t,\leq l)$-identifying codes. (OTD-codes for short), see [8] and [21], respectively.
Figure 1 illustrates examples of such codes in a small graph.
Figure 1: Minimum X-codes in a graph (the black vertices belong to the code), where (a) is an ID-code, (b) a ITD-code, (c) both an LD- and LTD-code, (d) both an OD- and OTD-code.
[…]
Problems of this type are intensively studied in the literature (see, for example, the internet bibliography containing over 500 articles around these topics maintained by Jean and Lobstein [15]) and have manifold applications, for example, in fault-detection in multiprocessor networks [16] or in locating threats/ intruders in facilities using sensor networks [24]. To model, for example, monitoring problems of the latter type, the rooms of a building are represented by vertices of a graph, doors between rooms as edges, and the task is to place monitoring devices at some vertices of the graph (a dominating or total-dominating set) such that the intruder (like a fire, a thief, or a saboteur) can be either detected directly in the room where the device is present or can be located with the help of a unique pattern of alerts sent out by other devices in some of the adjacent rooms. In the literature, several adaptations have been discussed to tackle further practical challenges, whereby some of the monitoring devices may become faulty over time:
  • Fault type 1: A monitoring device may lose the ability to distinguish the presence of an intruder in the room it is installed in or in an adjacent room.
  • Fault type 2: An intruder may disable the monitoring device in a room.
Monitoring without faulty or disabled devices leads to location [23], tackling faults of type 1 corresponds to closed-separation [16], whereas tackling faults of type 2 corresponds to open-separation [21, 14]. However, to the best of our knowledge, the literature still lacks the model of a monitoring system where faults of both type 1 and 2 may occur simultaneously. The purpose of this article is to introduce such a separation property.
To study various X-problems from a unifying point of view, reformulations of the above six X-problems in terms of covering problems in suitably constructed hypergraphs have been considered, for example, in [1, 2, 3, 8]. Given a graph $G=(V,E)$ and an X-problem, we look for a hypergraph $\mathcal{H}_{\textsc{X}}(G)=(V,\mathcal{F}_{\textsc{X}})$ so that $C\subseteq V$ is an X-code of $G$ if and only if $C$ is a cover of $\mathcal{H}_{\textsc{X}}(G)$, that is, a subset $C\subseteq V$ satisfying $C\cap F\neq\emptyset$ for all $F\in\mathcal{F}_{\textsc{X}}$. Then the covering number $\tau(\mathcal{H}_{\textsc{X}}(G))$, defined as the minimum cardinality of a cover of $\mathcal{H}_{\textsc{X}}(G)$, equals by construction the X-number $\gamma^{\rm{X}}(G)$. The hypergraph $\mathcal{H}_{\textsc{X}}(G)$ is called the X-hypergraph of $G$.
It is a simple observation that for an X-problem involving domination (respectively, total-domination), $\mathcal{F}_{\textsc{X}}$ needs to contain the closed (respectively, open) neighborhoods of all vertices of $G$. In order to encode the separation properties, that is, the fact that the intersection of an X-code with the neighborhood of each vertex is unique, it was suggested in [1, 2] to use the symmetric differences of the neighborhoods. Here, given two sets $A$ and $B$, their symmetric difference is defined by $A\bigtriangleup B=(A\setminus B)\cup(B\setminus A)$. In fact, it has been shown in [1, 2] that a code $C$ of a graph $G$ is
  • closed-separating if and only if $(N[u]\bigtriangleup N[v])\cap C\neq\emptyset$ for all distinct $u,v\in V$,
  • open-separating if and only if $(N(u)\bigtriangleup N(v))\cap C\neq\emptyset$ for all distinct $u,v\in V$,
  • locating if and only if $(N(u)\bigtriangleup N(v))\cap C\neq\emptyset$ for all $uv\in E$ and $(N[u]\bigtriangleup N[v])\cap C\neq\emptyset$ for all $uv\notin E$, where $u,v\in V$ are distinct.
Accordingly, for each of the already studied X-problems, the hyperedge set $\mathcal{F}_{\textsc{X}}$ of $\mathcal{H}_{\textsc{X}}(G)$ consists of a combination of exactly 3 different subsets as shown in Table 1, where
  • $N[G]$ (respectively, $N(G)$) denotes the set of all closed (respectively, open) neighborhoods of vertices in $G$,
  • $\bigtriangleup_{a}[G]$ (respectively, $\bigtriangleup_{a}(G)$) denotes the set of symmetric differences of closed (respectively, open) neighborhoods of all pairs of adjacent vertices in $G$,
  • $\bigtriangleup_{n}[G]$ (respectively, $\bigtriangleup_{n}(G)$) denotes the set of symmetric differences of closed (respectively, open) neighborhoods of all pairs of non-adjacent vertices in $G$.
X ID ITD LD LTD OD OTD
$N[G]$ $N(G)$ $N[G]$ $N(G)$ $N[G]$ $N(G)$
$\mathcal{F}_{\textsc{X}}$ $\bigtriangleup_{a}[G]$  $\bigtriangleup_{a}[G]$  $\bigtriangleup_{a}(G)$  $\bigtriangleup_{a}(G)$  $\bigtriangleup_{a}(G)$  $\bigtriangleup_{a}(G)$
 $\bigtriangleup_{n}[G]$  $\bigtriangleup_{n}[G]$  $\bigtriangleup_{n}[G]$  $\bigtriangleup_{n}[G]$  $\bigtriangleup_{n}(G)$  $\bigtriangleup_{n}(G)$
Table 1: The X-hypergraphs $\mathcal{H}_{\textsc{X}}(G)=(V,\mathcal{F}_{\textsc{X}})$ for the already considered X-problems, listing column-wise the 3 needed subsets of hyperedges
Analyzing this summary of hypergraph representations of the six already established X-problems, we observe that one separation property has not yet been considered in the literature, namely the one involving $\bigtriangleup_{a}[G]$ and $\bigtriangleup_{n}(G)$. The aim of this paper is to introduce the corresponding separation property, called full-separation, and to study it in combination with both domination and total-domination leading to full-separating-dominating codes (or FD-codes for short) and full-separating-total-dominating codes (or FTD-codes for short), respectively.
Note that not all graphs admit codes of all studied types. We accordingly address the conditions for the existence of FD- and FTD-codes, bounds for their size and their relation to codes of the other types in Section 2. In particular, we show that $\gamma^{\rm{FD}}(G)$ and $\gamma^{\rm{FTD}}(G)$ differ by at most one. Moreover, the problems of determining an X-code of minimum cardinality $\gamma^{\rm{X}}(G)$ in a graph $G$ have been shown to be NP-hard for all the previously studied X-problems [9, 10, 8, 21]. We show the same for FD- and FTD-codes in Section 3. Furthermore, we show that deciding whether the minimum FD- and FTD-codes of a graph have equal cardinality is NP-complete. This motivates to compare FD- and FTD-codes on different families of graphs as well as to study their relations to other X-codes in Section 4. We close with some concluding remarks and lines of future research.
2 Full-separation and related codes
In this section, we formally introduce the concept of full-separation, the related full-separating codes and address fundamental questions concerning the existence of these codes and bounds for their cardinalities.
Full-separation and related codes.
Let $G=(V,E)$ be a graph. A subset $C\subseteq V$ is a full-separating set of $G$ if, for each pair of distinct vertices $u,v\in V$, we have
$(N(v)\cap C)\setminus\{u\}\neq(N(u)\cap C)\setminus\{v\}.$
Moreover, a full-separating set $C\subseteq V$ is a full-separating-dominating code (FD-code for short) (respectively, a full-separating-total-dominating code (FTD-code for short)) if $C$ is also a dominating (respectively, total-dominating) set of $G$. The ITD-code from Figure 1 is also both a minimum FD- and FTD-code of the graph, Figure 2 illustrates further examples of such codes in a small graph.
Figure 2: Minimum X-codes in a graph (the black vertices belong to the code), where (a) is an FD-code, (b) an FTD-code.
In the following, we show some characterizations of full-separating sets.
Theorem 2.1 ($\star$)
Let $G=(V,E)$ be a graph. For a subset $C\subseteq V$, the following assertions are equivalent:
  1. (a)
    $C$ is full-separating.
  2. (b)
    $C$ has a non-empty intersection with $\big{(}N(u)\bigtriangleup N(v)\big{)}\setminus\{u,v\}$ for all distinct $u,v\in V$.
  3. (c)
    $C$ has a non-empty intersection with
    • $N[u]\bigtriangleup N[v]$ for all pairs of adjacent vertices $u,v\in V$, and
    • $N(u)\bigtriangleup N(v)$ for all pairs of non-adjacent vertices $u,v\in V$.
  4. (d)
    $N[u]\cap C\neq N[v]\cap C$ and $N(u)\cap C\neq N(v)\cap C$ for each pair of distinct vertices $u,v\in V$.
  5. (e)
    $C$ is both a closed-separating and an open-separating set of $G$.
Accordingly, by Theorem 2.1c, the hyperedge sets of both $\mathcal{H}_{\textsc{FD}}(G)$ and $\mathcal{H}_{\textsc{FTD}}(G)$ are composed of $N[G]$ for domination and $N(G)$ for total-domination, respectively, as well as $\bigtriangleup_{a}[G]$ and $\bigtriangleup_{n}(G)$ in both cases. Moreover, it turns out that a full-separating set incorporates both elements of closed-separation and open-separation. Thus, Theorem 2.1 also justifies how the separation property of full-separation models a monitoring system that tackles the faults of type 1 and 2 simultaneously.
Existence of FD- and FTD-codes.
It is known from the literature that the already studied X-codes may not exist in all graphs, see for example [8, 16, 13, 21]. Considering the existence of X-codes from the point of view as covers of the X-hypergraphs $\mathcal{H}_{\textsc{X}}(G)=(V,\mathcal{F}_{\textsc{X}})$, we immediately see that $G$ has no X-code if and only if $\mathcal{H}_{\textsc{X}}(G)$ has no cover if and only if $\emptyset\in\mathcal{F}_{\textsc{X}}$. In this context, it has been observed in [1, 2] that there is no X-code in a graph $G$ with $\mathcal{H}_{\textsc{X}}(G)$ involving
  • $N(G)$ if $G$ has isolated vertices, that is, vertices $v$ with $N(v)=\emptyset$;
  • $\bigtriangleup_{a}[G]$ if $G$ has closed twins, that is, adjacent vertices $u,v$ with $N[u]=N[v]$ (and thus $N[u]\bigtriangleup N[v]=\emptyset$);
  • $\bigtriangleup_{n}(G)$ if $G$ has open twins, that is, non-adjacent vertices $u,v$ with $N(u)=N(v)$ (and thus $N(u)\bigtriangleup N(v)=\emptyset$).
X ID ITD LD LTD FD FTD OD OTD
$N[G]$ $\mathbf{N(G)}$ $N[G]$ $\mathbf{N(G)}$ $N[G]$ $\mathbf{N(G)}$ $N[G]$ $\mathbf{N(G)}$
$\mathcal{F}_{\textsc{X}}$ $\mathbf{\Delta_{a}[G]}$  $\mathbf{\Delta_{a}[G]}$  $\bigtriangleup_{a}(G)$  $\bigtriangleup_{a}(G)$ $\mathbf{\Delta_{a}[G]}$  $\mathbf{\Delta_{a}[G]}$  $\bigtriangleup_{a}(G)$  $\bigtriangleup_{a}(G)$
 $\bigtriangleup_{n}[G]$  $\bigtriangleup_{n}[G]$  $\bigtriangleup_{n}[G]$  $\bigtriangleup_{n}[G]$  $\mathbf{\Delta_{n}(G)}$  $\mathbf{\Delta_{n}(G)}$  $\mathbf{\Delta_{n}(G)}$  $\mathbf{\Delta_{n}(G)}$
Table 2: The X-hypergraphs $\mathcal{H}_{\textsc{X}}(G)=(V,\mathcal{F}_{\textsc{X}})$, listing in boldface the subsets of hyperedges that may lead to non-existence of X-codes.
Table 2 illustrates which X-problems are concerned. Calling a graph $G$ X-admissible if $G$ has an X-code, we see, for example, from Table 2 that every graph $G$ is LD-admissible, whereas a graph $G$ is OTD-admissible if and only if $G$ has neither isolated vertices nor open twins. Moreover, we call a graph twin-free if it has neither closed nor open twins. Accordingly, we conclude the following regarding the existence of FD- and FTD-codes in graphs:
Corollary 1
A graph $G$ is
  • FD-admissible if and only if $G$ is twin-free;
  • FTD-admissible if and only if $G$ is twin-free and has no isolated vertex.
Since any two distinct isolated vertices of a graph are open twins with the empty set as both their open neighborhoods, Corollary 1 further implies the following.
Corollary 2
A FD-admissible graph has at most one isolated vertex.
Relations of X-numbers and resulting bounds on FD- and FTD-numbers.
Consider the set $\mathcal{H}^{*}(V)$ of all hypergraphs $\mathcal{H}=(V,\mathcal{F})$ on the same vertex set $V$. We define a relation $\prec$ on $\mathcal{H}^{*}(V)\times\mathcal{H}^{*}(V)$ by $\mathcal{H}\prec\mathcal{H}^{\prime}$ if, for every hyperedge $F^{\prime}$ of $\mathcal{H}^{\prime}$, there exists a hyperedge $F$ of $\mathcal{H}$ such that $F\subseteq F^{\prime}$. We can show:
Lemma 1 ($\star$)
If $\mathcal{H},\mathcal{H}^{\prime}\in\mathcal{H}^{*}(V)$ with $\mathcal{H}\prec\mathcal{H}^{\prime}$, then any cover of $\mathcal{H}$ is also a cover of $\mathcal{H}^{\prime}$. In particular, we have $\tau(\mathcal{H}^{\prime})\leq\tau(\mathcal{H})$.
Let Codes = {ID, ITD, LD, LTD, FD, FTD, OD, OTD}. It is easy to see that for every graph $G=(V,E)$, the X-hypergraphs $\mathcal{H}_{\textsc{X}}(G)=(V,\mathcal{F}_{\textsc{X}})$ belong to $\mathcal{H}^{*}(V)$ for all $X\in{\textsc{Codes}}$. Furthermore, it is clear that
  • $N[v]=N(v)\cup\{v\}$ for all vertices $v\in V$;
  • $N(u)\bigtriangleup N(v)=(N[u]\bigtriangleup N[v])\cup\{u,v\}$ for all adjacent vertices $u,v\in V$;
  • $N[u]\bigtriangleup N[v]=(N(u)\bigtriangleup N(v))\cup\{u,v\}$ for all non-adjacent vertices $u,v\in V$.
This observation combined with the above lemma implies:
Corollary 3
If for a graph $G=(V,E)$ and two problems $X,X^{\prime}\in{\textsc{Codes}}$, the hyperedges of $\mathcal{H}_{\textsc{X}}(G)=(V,\mathcal{F}_{\textsc{X}})$ and $\mathcal{H}_{{\textsc{X}}^{\prime}}(G)=(V,\mathcal{F}_{{\textsc{X}}^{\prime}})$ only differ in the
  • (a)
    neighborhoods such that $N(G)\subset\mathcal{F}_{\textsc{X}}$ and $N[G]\subset\mathcal{F}_{{\textsc{X}}^{\prime}}$, or
  • (b)
    symmetric differences of neighborhoods of adjacent vertices such that $\bigtriangleup_{a}[G]\\ \subset\mathcal{F}_{\textsc{X}}$ and $\bigtriangleup_{a}(G)\subset\mathcal{F}_{{\textsc{X}}^{\prime}}$, or
  • (c)
    symmetric differences of neighborhoods of non-adjacent vertices such that $\bigtriangleup_{n}(G)\subset\mathcal{F}_{\textsc{X}}$ and $\bigtriangleup_{n}[G]\subset\mathcal{F}_{{\textsc{X}}^{\prime}}$,
then we have $\mathcal{H}_{\textsc{X}}(G)\prec\mathcal{H}_{{\textsc{X}}^{\prime}}(G)$ in all the above cases which implies in particular $\gamma^{{\textsc{X}}^{\prime}}(G)\leq\gamma^{\textsc{X}}(G)$.
From Corollary 3 combined with the hypergraph representations from Table 2, we immediately obtain the relations between the X-numbers for all $X\in{\textsc{Codes}}$ shown in Figure 3.
Figure 3: The relations between the X-numbers for all $X\in{\textsc{Codes}}$, where $X^{\prime}\longrightarrow X$ stands for $\gamma^{{\textsc{X}}^{\prime}}(G)\leq\gamma^{\textsc{X}}(G)$ and solid (respectively, dotted and dashed) arrows refer to case (a) (respectively, (b) and (c)) of Corollary 3.
Note that most of the relations not involving FD- or FTD-numbers were already known before, see, for example, [2, 8]. However, from the global point of view on all X-numbers given in Figure 3, we see in particular that $\gamma^{\rm{LD}}(G)$ is a lower bound for all other X-numbers, whereas $\gamma^{\rm{FTD}}(G)$ is an upper bound for all other X-numbers in FTD-admissible graphs. Moreover, we can establish the following relation of FD- and FTD-numbers:
Theorem 2.2 ($\star$)
Let $G$ be an FD-admissible graph. If $G$ is a disjoint union of a graph $G^{\prime}$ and an isolated vertex, then we have $\gamma^{\rm{FD}}(G)=\gamma^{\rm{FTD}}(G^{\prime})+1$; otherwise, $\gamma^{\rm{FTD}}(G)-1\leq\gamma^{\rm{FD}}(G)\leq\gamma^{\rm{FTD}}(G)$ holds.
Combining the previous results, we further conclude the following lower bounds on FD- and FTD-numbers.
Corollary 4
Let $G$ be an FTD-admissible graph. We have
$\begin{array}[]{lcl}\gamma^{\rm{FTD}}(G)&\geq&\max(\gamma^{\rm{ITD}}(G),\gamma^{\rm{OTD}}(G),\gamma^{\rm{FD}}(G)),\\[2.84526pt] \gamma^{\rm{FD}}(G)&\geq&\max(\gamma^{\rm{ID}}(G),\gamma^{\rm{OD}}(G),\gamma^{\rm{LTD}}(G)-1,\gamma^{\rm{ITD}}(G)-1,\gamma^{\rm{OTD}}(G)-1).\\ \end{array}$
3 Complexity of the FD- and FTD-problems
In this section, we address the complexity of the FD- and FTD-problems. The decision versions of these two problems are formally stated as follows.
FD Input: $(G,k)$: A FD-admissible graph $G$ and a non-negative integer $k$. Question: Does there exist an FD-code $C$ of $G$ such that $|C|\leq k$?
FTD Input: $(G,k)$: A FTD-admissible graph $G$ and a non-negative integer $k$. Question: Does there exist an FTD-code $C$ of $G$ such that $|C|\leq k$?
Theorem 3.1
FTD is NP-complete.
Theorem 3.2
FD is NP-complete.
We also prove that, despite a difference of at most one as in Theorem 2.2, it is hard to decide if the FD- and FTD-numbers of a given input graph are the same or different. The decision version of this problem is defined as follows.
FD $\neq$ FTD Input: $(G,k)$: A FTD-admissible graph $G$ and a non-negative integer $k$. Question: Does there exist an FD-code $C^{\prime}$ and an FTD-code $C$ of $G$ such that $|C^{\prime}|<|C|\leq k$?
Theorem 3.3
FD $\neq$ FTD is NP-complete.
To prove the above three theorems, we use one single reduction from an instance $\psi$ of 3-SAT to an instance $(G^{\psi},k)$ of FD, FTD and FD $\neq$ FTD.
Reduction 1
The reduction takes as an input an instance $\psi=(A,B)$ of 3-SAT, where $A$ denotes the set of variables and $B$ denotes the set of clauses in $\psi$. Moreover, let there be $n=|A|$ variables and $m=|B|$ clauses in $\psi$. The reduction contructs a graph $G^{\psi}$ on $10n+3m$ vertices as follows (also refer to Figure 4):
  • For every variable $x\in A$, do the following (refer to Figure 4(a)):
    • Add 10 vertices named $v^{x}_{1}$, $v^{x}_{2}$, $v^{x}_{3}$, $w^{x}_{1}$, $w^{x}_{2}$, $s^{x}_{1}$, $s^{x}_{2}$, $s^{x}_{3}$, $z^{x}_{1}$ and $z^{x}_{2}$ (intuitively, the vertex $w^{x}_{1}$ will correspond to the literal $x$ and the vertex $w^{x}_{2}$ will correspond to the literal $\neg x$).
    • Add edges so that the graph induced by the vertices $v^{x}_{3}$, $w^{x}_{1}$, $w^{x}_{2}$, $s^{x}_{1}$, $s^{x}_{2}$ and $s^{x}_{3}$ is a $K_{6}$ minus the edges $w^{x}_{1}w^{x}_{2}$, $v^{x}_{1}s^{x}_{1}$, $v^{x}_{1}s^{x}_{2}$ and $v^{x}_{1}s^{x}_{3}$.
    • Add edges $v^{x}_{1}v^{x}_{2}$ and $v^{x}_{2}v^{x}_{3}$ making the vertices $v^{x}_{1}$, $v^{x}_{2}$ and $v^{x}_{3}$ induce a $P_{3}$; and add edges $s^{x}_{1}v^{z}_{1}$ and $s^{x}_{2}z^{x}_{2}$ making $z^{x}_{1}$ and $z^{x}_{2}$ pendant vertices and the vertices $s^{x}_{1}$ and $s^{x}_{2}$ their respective supports.
    Let the graph induced by the above 10 vertices be denoted as $G^{x}$ and be called the variable gadget corresponding the the variable $x\in A$.
  • For every clause $y\in B$, do the following (refer to Figure 4(b)):
    • Add 3 vertices named $u^{x}_{1}$, $u^{x}_{2}$ and $u^{x}_{3}$.
    • Add edges $u^{x}_{1}u^{x}_{2}$ and $u^{x}_{2}u^{x}_{3}$ thus making $u^{x}_{1}$, $u^{x}_{2}$ and $u^{x}_{3}$ induce a $P_{3}$.
    Let the $P_{3}$ induced by the above 3 vertices be denoted by $P^{y}$ and be called the clause gadget corresponding to the clause $y\in B$.
  • For all variables $x\in A$ and all clauses $y\in B$, if the literal $x$ is in a clause $y$, then add the edge $u^{y}_{1}w^{x}_{1}$; and if the literal $\neg x$ is in a clause $y$, then add the edge $u^{y}_{1}w^{x}_{2}$ (refer to Figure 4(c)).
$w^{x}_{1}$$w^{x}_{2}$$v^{x}_{1}$$v^{x}_{2}$$v^{x}_{3}$$s^{x}_{1}$$s^{x}_{2}$$s^{x}_{3}$$z^{x}_{1}$$z^{x}_{2}$
(a) The variable gadget $G^{x}$ corresponding to a variable $x\in A$.
$u^{y}_{1}$$u^{y}_{2}$$u^{y}_{3}$
(b) The clause gadget $P^{y}$ corresponding to a clause $y\in B$.
$w^{x}_{1}$$w^{x}_{2}$$v^{x}_{1}$$v^{x}_{2}$$v^{x}_{3}$$s^{x}_{1}$$s^{x}_{2}$$s^{x}_{3}$$z^{x}_{1}$$z^{x}_{2}$$u^{y}_{1}$$u^{y}_{2}$$u^{y}_{3}$
(c) The graph $G^{\psi}$ constructed from the instance $\psi$ of 3-SAT in Reduction 1.
Figure 4: Components of Reduction 1.
It can be verified that Reduction 1 is carried out in time polynomial in the inputs $m$ and $n$. The proofs of Theorems 3.1, Theorems 3.2 and 3.3 comprise of first proving the following lemmas. For the rest of this section, we fix the elements $\psi=(A,B)$, $|A|=n$, $|B|=m$ and the graph $G^{\psi}$ as defined so far.
Lemma 2 ($\star$)
If $\psi$ has a satisfying assignment, then there exists
  1. 1.
    an FTD-code $C$ of $G^{\psi}$ such that $|C|=7n+2m$; and
  2. 2.
    an FD-code $C^{\prime}$ of $G^{\psi}$ such that $|C^{\prime}|=7n+2m-1$.
Lemma 3 ($\star$)
The following statements are true.
  1. 1.
    If there exists an FTD-code $C$ of $G^{\psi}$ such that $|C|\leq 7n+2m$, then $\psi$ has a satisfying assignment.
  2. 2.
    If there exists an FD-code $C^{\prime}$ of $G^{\psi}$ such that $|C^{\prime}|\leq 7n+2m-1$, then $\psi$ has a satisfying assignment.
Proof (of Theorem 3.1)
The problem FTD clearly belongs to the class NP since it can be verified in polynomial-time if a given vertex subset of a graph $G$ is an FTD-code of $G$. To prove NP-completeness, we show that FTD is also NP-hard by showing that an instance $\psi=(A,B)$ with $|A|=n$ and $|B|=m$ is a yes-instance of 3-SAT if and only if $(G^{\psi},7n+2m)$ is a yes-instance of FTD, where the graph $G^{\psi}$ is as constructed in Reduction 1. In other words, we show that $\psi$ has a satisfying assignment if and only if there exists an FTD-code $C$ of $G^{\psi}$ such that $|C|\leq 7n+2m$. The last statement is true by Lemmas 2(1) and 3(1). ∎
Proof (of Theorem 3.2)
The problem FD clearly belongs to the class NP since it can be verified in polynomial-time if a given vertex subset of a graph $G$ is an FD-code of $G$. To prove NP-completeness, we show that FD is also NP-hard by showing that an instance $\psi=(A,B)$ with $|A|=n$ and $|B|=m$ is a yes-instance of 3-SAT if and only if $(G^{\psi},7n+2m-1)$ is a yes-instance of FTD, where the graph $G^{\psi}$ is as constructed in Reduction 1. In other words, we show that $\psi$ has a satisfying assignment if and only if there exists an FD-code $C$ of $G^{\psi}$ such that $|C|\leq 7n+2m-1$. The last statement is true by Lemmas 2(2) and 3(2). ∎
Proof (of Theorem 3.3)
The problem of FD $\neq$ FTD clearly belongs to the class NP since it can be verified in polynomial-time if given two vertex subsets $C$ and $C^{\prime}$ of a graph $G$ are FTD and FD-codes, respectively, of $G$. To prove NP-completeness, we also show that the problem is NP-hard by showing that an instance $\psi=(A,B)$ with $|A|=n$ and $|B|=m$ is a yes-instance of 3-SAT if and only if $(G^{\psi},7n+2m)$ is a yes-instance of FD $\neq$ FTD, where the graph $G^{\psi}$ is as constructed in Reduction 1. In other words, we show that $\psi$ has a satisfying assignment if and only if there exist a minimum FD-code $C^{\prime}$ and a minimum FTD-code $C$ of $G^{\psi}$ such that $|C^{\prime}|<|C|\leq 7n+2m$.
The necessary condition is true since, if $\psi$ has a satisfying assignment, then by Lemma 2, there exist an FD-code $C^{\prime}$ and an FTD-code $C$ of $G^{\psi}$ such that $|C^{\prime}|=7n+2m-1$ and $|C|=7n+2m$. To prove the sufficiency condition, let us assume that there exists an FD-code $C^{\prime}$ and an FTD-code $C$ of $G^{\psi}$ such that $|C^{\prime}|<|C|\leq 7n+2m$. Then, by Lemma 3(1), $\psi$ has a satisfying assignment. ∎
4 FD- and FTD-numbers of some graph families
In this section, we study the FD- and FTD-numbers of graphs belonging to some well-known graph families. In particular, Theorem 2.2 motivates us to compare the FD- and FTD-numbers for different families of graphs, whereas the relations to other X-numbers from Figure 3 raise questions about the incomparability of the FD-number with ITD-, OTD-, and LTD-numbers or the possible differences of FD- and FTD-numbers to their lower bounds in terms of other X-numbers due to Corollary 4. The latter questions motivate us to focus our study on families of twin-free graphs for which several other X-numbers have been already determined in the literature.
Paths and cycles.
In order to study the FD- and FTD-numbers of paths $P_{n}$ and cycles $C_{n}$, we first note that $P_{2}$ and $C_{3}$ have closed twins, $P_{3}$ and $C_{4}$ have open twins, but that all other paths and cycles are FD- and FTD-admissible. The proof of the following theorem giving the exact values of the FD- and FTD-numbers for paths and cycles is based on results by [21] and [6], respectively.
[…]
Figure 5: $P_{n}$ and $C_{n}$ (by joining the vertices $v_{1}$ and $v_{n}$ by an edge in each case). The set of black vertices represent an FTD-code.
Theorem 4.1 ($\star$)
Let $G$ be either a path $P_{n}$ for $n\geq 4$ or a cycle $C_{n}$ for $n\geq 5$. Moreover, let $n=6q+r$ for non-negative integers $q$ and $r\in[0,5]$. Then, we have
$\gamma^{\rm{FD}}(G)=\gamma^{\rm{FTD}}(G)=\begin{cases}4q+r,&\text{if $r\in[0,4]$};\\ 4q+4,&\text{if $r=5$}.\end{cases}$
Proof (sketch)
It can be shown that $\gamma^{\rm{FD}}(G)=\gamma^{\rm{FTD}}(G)$ when $G$ is either a path or a cycle of order at least $4$ and $5$, respectively. Hence, it is enough to prove the result for $\gamma^{\rm{FTD}}(G)$ only. First, we prove that $\gamma^{\rm{FTD}}(G)\leq 4q+r$ if $r\in[4]$ and $\gamma^{\rm{FTD}}(G)\leq 4q+4$ if $r=5$ by devising a linear-time algorithm to build an FTD-code of $G$ (as shown in Figure 5) of cardinality $4q+r$ if $r\in[4]$ and $44q+$ 4 if $r=5$, respectively. Then for $G\cong P_{n}$, the result follows immediately by the fact that $\gamma^{\rm{FTD}}(P_{n})\geq\gamma^{\rm{OTD}}(P_{n})$ and the result in [21] that $\gamma^{\rm{OTD}}(P_{n})=4q+r$ if $r\in[4]$ and $44q+$ 4 if $r=5$. Moreover, for $G\cong C_{n}$, by using a result from [6], it can be shown that we only need to prove $\gamma^{\rm{FTD}}(C_{n})\geq 4q+3$ whenever $n=6q+3$. We assume on the contrary that there exists a minimum FTD-code $C$ of $G$ such that $|C|\leq 4q+2$. Now, it can be verified that, for every vertex $v$ of $G$, there exists a set $S_{v}\subset C$ such that $v\in S_{v}$, $|S_{4}|\geq 4$ and $G[S_{v}]\cong P_{|S_{v}|}$. With our assumption that $|C|\leq 4q+2$, it can be shown that there exists one vertex $w$ of $G$ for which $|S_{w}|\geq 5$. Then by contracting one edge in $G[S_{w}]\cong P_{|S_{w}|}$ we find a cycle $C_{6q+2}$ and an FTD-code of cardinality at most $14q+$ 1 thus contradicting the result in [6] that $\gamma^{\rm{FTD}}(C_{6q+2})=4q+2$. This proves the result. ∎
Half-graphs.
A graph $G=(U\cup W,E)$ is bipartite if its vertex set can be partitioned into two stable sets $U$ and $W$ so that every edge of $G$ has one endpoint in $U$ and the other in $W$. For any integer $k\geq 1$, the half-graph $B_{k}=(U\cup W,E)$ is the bipartite graph with vertices in $U=\{u_{1},\ldots,u_{k}\}$, $W=\{w_{1},\ldots,w_{k}\}$ and edges $u_{i}w_{j}$ if and only if $i\leq j$ (see Figure 6(a)). In particular, we have $B_{1}=K_{2}$ and $B_{2}=P_{4}$. Moreover, we clearly see that half-graphs are connected and twin-free for all $k\geq 2$ and hence, are both FD- and FTD-admissible in this case.
In [11] it was shown that the only graphs whose OTD-numbers equal the order of the graph are the disjoint unions of half-graphs. In particular, we have $\gamma^{\rm{OTD}}(B_{k})=2k$. Combining this result with Corollary 3(b) showing that $\gamma^{\rm{OTD}}(G)\leq\gamma^{\rm{FTD}}(G)$ holds for all graphs and Theorem 2.2 showing $\gamma^{\rm{FD}}(G+K_{1})=\gamma^{\rm{FTD}}(G)+1$ together yields:
Corollary 5
For a graph $G=(V,E)$ of order $n$ being the disjoint union of half-graphs, we have $\gamma^{\rm{FTD}}(G)=n$ and $\gamma^{\rm{FD}}(G+K_{1})=\gamma^{\rm{FTD}}(G)+1=n+1$.
Hence, there are graphs where the FD- and FTD-numbers equal the order of the graph. A further example is $B_{2}=P_{4}$ with $\gamma^{\rm{FD}}(P_{4})=4$. On the other hand, we can prove:
Theorem 4.2 ($\star$)
For a half-graph $B_{k}$ with $k\geq 3$, we have $\gamma^{\rm{FD}}(B_{k})=2k-1$.
Hence, for half-graphs $B_{k}$ with $k\geq 3$, FD- and FTD-numbers differ. This result combined with Theorem 2.2 shows that $G=B_{k}+K_{1}$ for $k\geq 3$ are examples of graphs where $\gamma^{\rm{FD}}(G)$ is larger than the sum of the FD-numbers of its components. Moreover, half-graphs are extremal graphs for the lower bounds $\gamma^{\rm{OTD}}(G)\leq\gamma^{\rm{FTD}}(G)$ and $\gamma^{\rm{OD}}(G),\gamma^{\rm{OTD}}(G)-1\leq\gamma^{\rm{FD}}(G)$ (the latter by combining Theorem 4.2 with results from [8] and [11] on OD- and OTD-numbers of half-graphs, respectively).
(a) The half-graph $B_{6}$.
(b) $P_{4}$
(c) $H_{3}$: net
(d) $\overline{H}_{3}$: sun
Figure 6: The black vertices in the figures represent minimum FD-codes. For Figures 6(b), 6(c) and 6(d), they also represent minimum FTD-codes. The grey vertex in Figures 6(a) depicts an extra vertex required to complete an FD- to the FTD-code.
Headless spiders.
[…]
A headless spider is a split graph with $Q=\{q_{1},\ldots,q_{k}\}$ and $S=\{s_{1},\ldots,s_{k}\}$. In addition, a headless spider is thin (respectively, thick) if $s_{i}$ is adjacent to $q_{j}$ if and only if $i=j$ (respectively, $i\neq j$). By definition, it is clear that the complement of a thin headless spider $H_{k}$ is a thick headless spider $\overline{H}_{k}$, and vice-versa. We have, for example, $H_{2}=\overline{H}_{2}=P_{4}$ and the two headless spiders $H_{3}=net$ and $\overline{H}_{3}=sun$ are depicted in Figures 6(b), 6(c) and 6(d), respectively. Moreover, it is easy to check that thin and thick headless spiders have no twins.
[…]
Theorem 4.3 ($\star$)
For a thin headless spider $H_{k}=(Q\cup S,E)$ with $k\geq 4$, we have $\gamma^{\rm{FD}}(H_{k})=2k-2$ and $\gamma^{\rm{FTD}}(H_{k})=2k-1$.
It is known from [2, 8] that $\gamma^{\rm{LD}}(\overline{H}_{k})=\gamma^{\rm{LTD}}(\overline{H}_{k})=k-1$ and $\gamma^{\rm{OD}}(\overline{H}_{k})=\gamma^{\rm{OTD}}(\overline{H}_{k})=k+1$ holds for all $k\geq 3$. Furthermore, $\gamma^{\rm{ID}}(\overline{H}_{k})=k$ for $k\geq 3$ was shown in [3]. We next determine the X-numbers of thick headless spiders for $X\in\{FD,FTD,ITD\}$.
Theorem 4.4 ($\star$)
For a thick headless spider $\overline{H}_{k}=(Q\cup S,E)$ with $k\geq 4$, we have $\gamma^{\rm{ITD}}(\overline{H}_{k})=k+1$ and $\gamma^{\rm{FD}}(\overline{H}_{k})=\gamma^{\rm{FTD}}(\overline{H}_{k})=2k-2$.
Hence, for thick headless spiders, FD- and FTD-numbers are equal. This shows that thick headless spiders are further extremal cases for the lower bound $\gamma^{\rm{FD}}(G)$ $\leq\gamma^{\rm{FTD}}(G)$, whereas thin headless spiders are extremal cases for the lower bounds $\gamma^{\rm{ITD}}(G)\leq\gamma^{\rm{FTD}}(G)$ and $\gamma^{\rm{ITD}}(G)-1\leq\gamma^{\rm{FD}}(G)$.
5 Concluding remarks
In this paper, we studied problems in the domain of identification problems, where the objective is to distinguish any two vertices of a graph by their unique neighborhoods in a suitably chosen dominating or total-dominating set. To study such problems under a unifying point of view, reformulations of the already studied problems in terms of covering problems in suitably constructed hypergraphs have been suggested in [1, 2, 3, 8], where the hypergraphs are composed of different combinations of closed or open neighborhoods and their symmetric differences (Table 1). Analyzing these hypergraph representations, we observed that one separation property has not yet been studied, namely the one involving $\bigtriangleup_{a}[G]$ and $\bigtriangleup_{n}(G)$.
The contribution of this paper is to introduce the corresponding separation property, called full-separation, and to study it in combination with both domination and total-domination. We characterize full-separating sets in different ways to show that they combine the requirements of both closed- and open separation. This is also reflected in the conditions for the existence of FD- and FTD-codes in graphs (Corollary 1). Based on the hypergraph representations, we devise relations between all X-numbers (Figure 3) and particularly between the FD- and FTD-numbers of a graph, showing that they differ by at most one (Theorem 2.2). That way, we detect that the FTD-number is an upper bound for all other X-numbers and deduce lower bounds for the FD- and FTD-numbers of a graph in terms of the other X-numbers (Corollary 4), again in accordance with their property requiring both closed- and open-separation. As is the case for the previously studied X-numbers, we further showed that deciding whether the FD- or FTD-number of a graph is below a given value is NP-complete (Theorems 3.1 and 3.2). Moreover, we show that, despite a difference of at most one between the FD- and FTD-numbers, it is in general NP-complete to decide if the two numbers differ on a graph. The latter result motivates us to compare the FD- and FTD-numbers on several graphs, providing graph families where they are equal (paths $P_{n}$ with $k\geq 4$, cycles $C_{n}$ with $k\geq 5$, thick headless spiders $\overline{H}_{k}$ with $k\geq 4$) or where they differ (half-graphs $B_{k}$ with $k\geq 3$, thin headless spiders $H_{k}$ with $k\geq 4$).
Moreover, comparing the FD- or FTD-numbers on the studied graph families with the other X-numbers known from the literature, we deduce that the FD-number is in general incomparable with the OTD-number. That is to say that for some graph families, for example half-graphs $G$, we have $\gamma^{\rm{FD}}(G)<\gamma^{\rm{OTD}}(G)$, and for other graph families, like thin headless spiders $G$, we have $\gamma^{\rm{OTD}}(G)<\gamma^{\rm{FD}}(G)$. It turns out that the FD-number is similarly incomparable with the ITD-number (considering their values for thin and thick headless spiders). However, we could not yet establish such an incomparability with the LTD-number. On the other hand, we note here that other linear comparisons between these numbers can possibly be established. For example, we believe that upper bounds of the form
$\gamma^{\rm{FTD}}(G)$ $\leq\min(2\gamma^{\rm{OTD}}(G)+\alpha,2\gamma^{\rm{ITD}}(G)+\beta);\text{ and}$
$\gamma^{\rm{FD}}(G)$ $\leq\min(2\gamma^{\rm{OD}}(G)+\alpha,2\gamma^{\rm{ID}}(G)+\beta),$
where $\alpha$ and $\beta$ are some integer constants, can be obtained by methods similar to those used in [22] to, for instance, show that $\gamma^{\rm{ID}}(G)\leq 2\gamma^{\rm{OTD}}(G)$. Finally, we could also provide examples of extremal cases for most of the lower bounds on FD- FTD-numbers from Corollary 4.
Our lines of future research include to pursue these studies of the FD- or FTD-codes by involving more graph families, for example, to find more examples of graphs where the FD- or FTD-numbers are not close to the order of the graph or the FD-number is smaller than the LTD-number (which would show that also FD- and LTD-numbers are in general incomparable). In general it would be interesting to identify or even characterize the extremal cases, that is, graphs for which the FD- or FTD-numbers equal the upper bound (the order of the graph) or one of its lower bounds in terms of other X-numbers.
Finally, it would be interesting to study a broader set of problem variants in the area of domination / separation from a hypergraph point of view towards a more general classification of such problems. In this context, it is clear that the hypergraphs encoding domination (respectively, total-domination) are only composed of closed (respectively, open) neighborhoods, whereas hypergraphs encoding only the separation properties only contain the corresponding symmetric differences of neighborhoods for all pairs of vertices. Here, it would be interesting to see if other restrictions (for example, only to $\bigtriangleup_{a}[G]$ or only to $N(G)$ and $\bigtriangleup_{a}(G)$) encode further studied problems from the literature and then similar relations as in Figure 3 can be deduced to a large set of graph parameters.
Appendix
In the appendix, we present the proofs of the results that have been omitted due to restrictions to the page limit.
Further terminologies and notations.
A vertex $w\in V$ is called full-separation forced if $\{w\}=(N(u)\setminus\{v\})\bigtriangleup(N(v)\setminus\{u\})$ for some distinct pair of vertices $u,v\in V$.
Observation A.1
Let $G=(V,E)$ be a graph and $u,v$ be any two distinct vertices of $G$. Then, we have
$(N(u)\bigtriangleup N(v))\setminus\{u,v\}=\left\{\begin{array}[]{lll}N[u]\bigtriangleup N[v]&\subset N(u)\bigtriangleup N(v),&\text{ if $uv\in E$};\\ N(u)\bigtriangleup N(v)&\subset N[u]\bigtriangleup N[v],&\text{ if $uv\notin E$.}\end{array}\right.$
In particular, therefore, the set $(N(u)\bigtriangleup N(v))\setminus\{u,v\}$ is a subset of both $N(u)\bigtriangleup N(v)$ and $N[u]\bigtriangleup N[v]$ for all distinct $u,v\in V$.
2 Proofs omitted in Section 2: Full-separation and related codes
Theorem 2.1
Let $G=(V,E)$ be a graph. For a subset $C\subseteq V$, the following assertions are equivalent:
  1. (a)
    $C$ is full-separating.
  2. (b)
    $C$ has a non-empty intersection with $\big{(}N(u)\bigtriangleup N(v)\big{)}\setminus\{u,v\}$ for all distinct $u,v\in V$.
  3. (c)
    $C$ has a non-empty intersection with
    • $N[u]\bigtriangleup N[v]$ for all pairs of adjacent vertices $u,v\in V$, and
    • $N(u)\bigtriangleup N(v)$ for all pairs of non-adjacent vertices $u,v\in V$.
  4. (d)
    $N[u]\cap C\neq N[v]\cap C$ and $N(u)\cap C\neq N(v)\cap C$ for each pair of distinct vertices $u,v\in V$.
  5. (e)
    $C$ is both a closed-separating and an open-separating set of $G$.
Proof
Let $u$ and $v$ be two distinct vertices of $G$.
a $\iff$ b: The result follows from the fact that, for any set $S\subset V$,
$(N(u)\cap S)\setminus\{v\}\neq(N(v)\cap S)\setminus\{u\}$
$\iff$ $\big{(}(N(u)\setminus\{v\})\bigtriangleup(N(v)\setminus\{u\})\big{)}\cap S\neq\emptyset$
$\iff$ $\big{(}(N(u)\bigtriangleup N(v))\setminus\{u,v\}\big{)}\cap S\neq\emptyset.$
b $\iff$ c: The result follows by Observation A.1.
b $\implies$ d: The result again follows by Observation A.1.
d $\implies$ b: First, let $u$ and $v$ be adjacent in $G$. Then, we have $(N[u]\bigtriangleup N[v])\cap C\neq\emptyset$. Therefore, Observation A.1 implies that $C$ has non-empty intersection with $\big{(}N(u)\bigtriangleup N(v)\big{)}\setminus\{u,v\}$. We now assume that $u$ and $v$ are non-adjacent in $G$. Then, we have $(N(u)\bigtriangleup N(v))\cap C\neq\emptyset$. Therefore, again Observation A.1 implies that $C$ has non-empty intersection with $\big{(}N(u)\bigtriangleup N(v)\big{)}\setminus\{u,v\}$. This proves the result.
d $\iff$ e: The result follows by definitions of closed- and open-separating sets. ∎
Relations of X-numbers and resulting bounds on FD- and FTD-numbers.
Lemma 1
If $\mathcal{H},\mathcal{H}^{\prime}\in\mathcal{H}^{*}(V)$ with $\mathcal{H}\prec\mathcal{H}^{\prime}$, then any cover of $\mathcal{H}$ is also a cover of $\mathcal{H}^{\prime}$. In particular, we have $\tau(\mathcal{H}^{\prime})\leq\tau(\mathcal{H})$.
Proof
Let $\mathcal{H},\mathcal{H}^{\prime}\in\mathcal{H}^{*}(V)$ with $\mathcal{H}=(V,\mathcal{F})$, $\mathcal{H}^{\prime}=(V,\mathcal{F}^{\prime})$ and assume that $\mathcal{H}\prec\mathcal{H}^{\prime}$. Moreover, let $C\subseteq V$ of $\mathcal{H}$ be a cover of $\mathcal{H}$. Now, let $F^{\prime}\in\mathcal{F}^{\prime}$ be any hyperedge. Since $\mathcal{H}\prec\mathcal{H}^{\prime}$, there exists a hyperedge $F$ of $\mathcal{H}$ such that $F\subseteq F^{\prime}$. Then $C\cap F\neq\emptyset$, as $C$ is a cover of $\mathcal{H}$. This implies that $C\cap F^{\prime}\neq\emptyset$ as well. Hence, $C$ interesects every hyperedge of $\mathcal{H}^{\prime}$ and thus, is also a cover of $\mathcal{H}^{\prime}$. The second statement follows by the fact that, if $C$ is minimum cover of $\mathcal{H}$, then $C$ is also a cover of $\mathcal{H}^{\prime}$ and hence, $\tau(\mathcal{H}^{\prime})\leq|C|=\tau(\mathcal{H})$. ∎
Observation A.2
The following are true for a twin-free graph $G=(V,E)$.
  1. (a)
    If $G$ is isolate-free, then every support vertex of $G$ belongs to a total-dominating set of $G$.
  2. (b)
    If $w$ is a full-separation forced vertex of $G$, then $w\in C$ for any full-separating set $C$ of $G$.
  3. (c)
    If $G$ is FD-admissible and $C$ is an FD-code of $G$, then there exists at most one vertex $v$ of $G$ such that $N(v)\cap C=\emptyset$.
  4. (d)
    If $G$ is FTD-admissible, then an FTD-code of $G$ is an X-code of $G$ for all X $\in$ Codes. In particular, $\gamma^{\rm{X}}(G)\leq\gamma^{\rm{FTD}}(G)$.
  5. (e)
    Let $G$ be FTD-admissible such that $\gamma^{\rm{FD}}(G)\neq\gamma^{\rm{FTD}}(G)$, then any minimum FD-code is not a total-dominating set of $G$. In particular, for every minimum FD-code $C$ of $G$, there exists exactly one vertex $v_{C}\in V$ such that $N(v_{C})\cap C=\emptyset$.
  6. (f)
    Let $G$ be FTD-admissible. Then, a set $C\subset V$ is an FTD-code of $G$ if and only if $C$ is a total-dominating set of $G$ and full-separates any distinct $u,v\in V$ such that $d(u,v)\leq 2$.
  7. (g)
    Let $G$ be FD-admissible. Then, a set $C\subset V$ is an FD-code of $G$ if and only if
    1. (i)
      $C$ is a dominating set of $G$ with at most one vertex $w\in V$ such that $N(w)\cap C=\emptyset$; and
    2. (ii)
      $C$ full-separates any distinct $u,v\in V$ such that $d(u,v)\leq 2$.
Proof
The following are the proofs of the observations.
  1. (a)
    For a support vertex $s\in V$ and its adjacent pendant vertex $v\in V$, since $N(v)=\{s\}$, the result follows directly from the definition of a total-dominating set of $G$.
  2. (b)
    This result too follows directly from the definitions of a full-separating set and a full-separation forced vertex.
  3. (c)
    If there exist two distinct vertices $u,v\in V$ such that $N(u)\cap C=\emptyset=N(v)\cap C$, then by definition, $C$ is not a full-separating set of $G$, a contradiction. Hence, the result is true.
  4. (d)
    The result is true by the successive use of the statements (a), (b) and (c) in Corollary 3 (or equivalently, by following the arrows in Figure 3).
  5. (e)
    Let $C$ be a minimum FD-code of $G$. On the contrary, if $C$ is a total-dominating set of $G$, then $C$ is also an FTD-code of $G$. Therefore, $\gamma^{\rm{FTD}}(G)\leq|C|=\gamma^{\rm{FD}}(G)$. Moreover, Observation A.2d implies that we have $\gamma^{\rm{FD}}(G)\leq\gamma^{\rm{FTD}}(G)$. Therefore, we have $\gamma^{\rm{FD}}(G)=\gamma^{\rm{FTD}}(G)$, a contradiction and hence, the result follows.
  6. (f)
    The necessary condition follows by the definition of full-separation. To prove the sufficient condition, it is enough to show that any two vertices $u,v\in V$ with $d(u,v)\geq 3$ are full-separated by $C$, that is, open-separated by $C$ (since $uv\notin E$). Now, $d(u,v)\geq 3$ implies that $N(u)\bigtriangleup N(v)=N(u)\cup N(v)$. Since $C$ is also a total-dominating set of $G$, we have $N(u)\cap C\neq\emptyset$ implying that $(N(u)\bigtriangleup N(v))\cap C\neq\emptyset$ as well. Hence, the pair $u,v$ are full-separated by $C$.
  7. (g)
    To prove the necessity part of the statement, we see that condition (i) follows from Observation A.2c and condition (ii) follows directly from the definition of full-separation. Therefore, to prove the sufficiency part of the statement, we assume that $C$ satisfies both conditions (i) and (ii). Then, it is enough to show that $C$ full-separates every pair $u,v$ of distinct vertices of $G$ such that $d(u,v)\geq 3$, that is, open-separates $u,v$ (since $uv\notin E$). Therefore, for two such vertices $u,v\in V$ with $d(u,v)\geq 3$, we again have $N(u)\triangle N(v)=N(u)\cup N(v)$. If $C$ does not open-separate $u$ and $v$, then it implies that $N(u)\cap C=N(v)\cap C=\emptyset$ thus contradicting condition (i). Therefore, $C$ full-separates $u$ and $v$ and this proves the result. ∎
Theorem 2.2
Let $G$ be an FD-admissible graph. If $G$ is a disjoint union of a graph $G^{\prime}$ and an isolated vertex, then we have $\gamma^{\rm{FD}}(G)=\gamma^{\rm{FTD}}(G^{\prime})+1$; otherwise, $\gamma^{\rm{FTD}}(G)-1\leq\gamma^{\rm{FD}}(G)\leq\gamma^{\rm{FTD}}(G)$ holds.
Proof
Let us first assume that $G$ is a disjoint union of a graph $G^{\prime}$ and an isolated vertex $v$. Moreover, let $C$ be a minimum FD-code of $G$. Then, $v\in C$ in order for $C$ to dominate $v$. This implies that $N(v)\cap C=\emptyset$. Therefore, by Observation A.2c, the set $C^{\prime}=C\setminus\{v\}$ must total-dominate all vertices in $V\setminus\{v\}$. Moreover, as $v$ is an isolated vertex of $G$, the set $C^{\prime}$ is also a full-separating set of $G^{\prime}$ and hence, an FTD-code of $G^{\prime}$. This implies that $\gamma^{\rm{FTD}}(G^{\prime})\leq|C|-1=\gamma^{\rm{FD}}(G)-1$, that is, $\gamma^{\rm{FD}}(G)\geq\gamma^{\rm{FTD}}(G^{\prime})+1$. On the other hand, if $C^{\prime}$ is an FTD-code of $G^{\prime}$, then $C^{\prime}\cup\{v\}$ is an FD-code of $G$. This implies that $\gamma^{\rm{FD}}(G)\leq|C^{\prime}|+1=\gamma^{\rm{FTD}}(G^{\prime})+1$ and hence, the result follows.
Let us now assume that the graph $G$ has no isolated vertices. Then, $G$ is also FTD-admissible. Now, the inequality $\gamma^{\rm{FD}}(G)\leq\gamma^{\rm{FTD}}(G)$ holds by Observation A.2d. To prove the other inquality, we let $C$ be a minimum FD-code of $G$. Then, we have $|C|=\gamma^{\rm{FD}}(G)$. Now, if $C$ is also an FTD-code of $G$, then the result holds trivially. Therefore, let us assume that $C$ is not an FTD-code of $G$, that is, in particular, $\gamma^{\rm{FD}}(G)\neq\gamma^{\rm{FTD}}(G)$. Therefore, by Observation A.2e, there exists a vertex $v_{C}$ of $G$ such that $N(v_{C})\cap C=\emptyset$. Moreover, by Observation A.2c, $C\cap N(v)=\emptyset$ for any $v\in V$ if and only if $v=v_{C}$. Since $G$ has no isolated vertices, the vertex $v_{C}$ has a neighbor, say $u$, in $G$. Then, $C\cup\{u\}$ is a total-dominating set of $G$ and is also a full-separating set of $G$. Therefore, we have $\gamma^{\rm{FTD}}(G)\leq|C|+1=\gamma^{\rm{FD}}(G)+1$. ∎
3 Proofs omitted in Section 4: Complexities of the FD- and FTD-problems
Lemma A.1
Let $S$ be a full-separating set of $G^{\psi}$.
  1. (a)
    If $S$ is either an FD- or an FTD-code of $G^{\psi}$, then for all $y\in B$, we have $|V(P^{y})\cap S|\geq 2$.
  2. (b)
    If $S$ is an FTD-code, then we have $|(V(G^{x})\setminus\{w^{x}_{1},w^{x}_{2}\})\cap S|\geq 6$ for all $x\in A$.
  3. (c)
    If $S$ is an FD-code, then except possibly for one $x^{\prime}\in A$ for which $|(V(G^{x^{\prime}})\setminus\{w^{x^{\prime}}_{1},w^{x^{\prime}}_{2}\})\cap S|\geq 5$, we have $|(V(G^{x})\setminus\{w^{x}_{1},w^{x}_{2}\})\cap S|\geq 6$ for all $x\in A\setminus\{x^{\prime}\}$.
Proof
(a): Let $S$ be either an FD- or an FTD-code of $G^{\psi}$ and let $y\in B$. Since $u^{y}_{3}$ is a pendant vertex of $G^{\psi}$, we must have |$|\{u^{y}_{2},u^{y}_{3}\}\cap S|\geq 1$. Moreover, the vertex $u^{x}_{1}$ is full-separation forced with respect to the pair $u^{x}_{2},u^{x}_{3}$. Therefore, by Observation A.2b, we also have $u^{x}_{1}\in S$. This proves (a).
[…]
This proves parts (b) and (c). ∎
Lemma 2
If $\psi$ has a satisfying assignment, then there exists
  1. 1.
    an FTD-code $C$ of $G^{\psi}$ such that $|C|=7n+2m$; and
  2. 2.
    an FD-code $C^{\prime}$ of $G^{\psi}$ such that $|C^{\prime}|=7n+2m-1$.
Proof
First, we show that there exists an FTD-code $C$ of $G^{\psi}$ such that $|C|=7n+2m$. To that end, we build a vertex subset $C$ of $G^{\psi}$ as follows.
  • For all $x\in A$ and $y\in B$, pick the vertices $u^{y}_{1},u^{y}_{2},v^{x}_{1},v^{x}_{2},z^{x}_{1},z^{x}_{2},s^{x}_{1},s^{x}_{2}$ in $C$.
  • For any variable $x\in A$, either $x$ or $\neg x$ belongs to a clause $y\in B$ such that the variable is assigned 1 in the satisfying assignment of $\psi$. If $x$ belongs to $y$, then pick $w^{x}_{1}$ in $C$; and if $\neg x$ belongs to $y$, then pick $w^{x}_{2}$ in $C$.
Note that, by construction, we have $|C|=7n+2m$. We now show that $C$ is an FTD-code of $G^{\psi}$. Notice that, again by construction, the set $C$ is a total-dominating set of $G^{\psi}$. Therefore, by Observations A.2f, it is enough to show that $C$ full-separates all pairs of distinct vertices of $G^{\psi}$ with distance at most 2 between them. For every $x\in A$ and $y\in B$, the set $C$ full-separates
  • the vertices $v^{x}_{2},v^{x}_{3}$ by $v^{x}_{1}$; the vertices $u^{y}_{2},u^{y}_{3}$ by $u^{y}_{1}$; and the vertices $v^{x}_{1},u^{y}_{1}$ by $u^{y}_{2}$.
  • It full-separates the vertices $v^{x}_{1},v^{x}_{3}$, by either $w^{x}_{1}$ or $w^{x}_{2}$ according to whether $(x,\neg x)=(1,0)$ or $(x,\neg x)=(0,1)$, respectively.
  • It full-separates the vertices $u^{y}_{1},u^{y}_{3}$ by either $w^{x}_{1}$ or $w^{x^{\prime}}_{2}$ for some $x,x^{\prime}\in A$ according to whether $x$ (assigned to 1) belongs to the clause $y$; or $\neg x^{\prime}$ (assigned to 1) belongs to $y$.
  • It full-separates the vertices $w^{x}_{1},w^{x}_{2}$ by $u^{y}_{1}$ for some clause $y\in B$ (to which exactly one of $w^{x}_{1}$ and $w^{x}_{2}$ belongs)
  • For any $i\in[2]$ and $j\in[3]$, the set $C$ full-separates the vertices $w^{x}_{i},s^{x}_{j}$ by $v^{x}_{1}$.
  • For any $i,j\in[2]$, the set $C$ full-separates the vertices $w^{x}_{i},z^{x}_{j}$ by $v^{x}_{1}$; the vertices $w^{x}_{i},u^{y}_{j}$ by $v^{x}_{1}$; and the vertices $w^{x}_{i},v^{y}_{j}$ by $s^{x}_{2}$.
  • For any $i,j\in[3]$, it full-separates the vertices $s^{x}_{i},s^{x}_{j}$, by either $z^{x}_{1}$ or $z^{x}_{2}$.
  • For any $i\in[3]$, it full-separates the vertices $s^{x}_{i},v^{x}_{1}$, by $z^{x}_{i}$; and the vertices $s^{x}_{i},u^{y}_{1}$, by $z^{x}_{i}$.
  • $C$ full-separates the vertices $s^{x}_{i},z^{x}_{j}$, by either $w^{x}_{1}$ or $w^{x}_{2}$ according to whether $(x,\neg x)=(1,0)$ or $(x,\neg x)=(0,1)$, respectively, where $i,j\in[2]$;
Finally, the only vertices in different variable gadgets (respectively, clause gadgets) with distance 2 between them are of the form $w^{x}_{i}\in G^{x}$ and $w^{x^{\prime}}_{j}\in G^{x^{\prime}}$ (respectively, $u^{y}_{1}\in P^{y}$ and $u^{y^{\prime}}_{1}\in P^{y^{\prime}}$), where $x,x^{\prime}\in A$ (respectively, $y,y^{\prime}\in B$) are distinct. However, the set $C$ full-separates any such pair $w^{x}_{i},w^{x^{\prime}}_{j}$ by $v^{x}_{1}$ and the pair $u^{y}_{1},u^{y^{\prime}}_{1}$ by $u^{y}_{2}$. This proves that $C$ is a full-separating sets of $G^{\psi}$.
We now let $C^{\prime}=C\setminus\{s^{x^{\prime}}_{1}\}$, where $x^{\prime}$ is a fixed variable in $A$. Therefore, we have $|C^{\prime}|=7n+2m-1$. Hence, it is enough to show that $C^{\prime}$ is an FD-code of $G^{\psi}$. First of all, it can be verified that $C^{\prime}$ is a dominating set of $G^{\psi}$. We now show that $C^{\prime}$ is also a full-separating set of $G^{\psi}$. To do so, notice that $N(z^{x^{\prime}}_{1})\cap C^{\prime}=\emptyset$ for only the variable $x^{\prime}\in A$. Therefore, by Observations A.2g, it is enough to show that $C^{\prime}$ full-separates all pairs of distinct vertices of $G^{\psi}$ with distance at most 2 between them. In other words, we consider for full-separation by $C^{\prime}$ the exact same pairs of vertices we had considered for full-separation by $C$ in the previous paragraph. Moreover, we notice that in the previous paragraph, $C$ does not full-separate any pair of vertices of $G^{\psi}$ by the vertex $s^{x}_{1}$ for any $x\in A$. This implies that each of those pairs of vertices of $G^{\psi}$ is also full-separated by $C^{\prime}$ as well and hence, this proves that $C^{\prime}$ too is a full-separating set of $G^{\psi}$. ∎
Lemma A.2
If there exists a full-separating set $S$ of $G^{\psi}$ such that |$|\{w^{x}_{1},w^{x}_{2}\}\cap S|=1$, then $\psi$ has a satisfying assignment.
Proof
Let $S$ be a full-separating set of $G^{\psi}$ such that |$|\{w^{x}_{1},w^{x}_{2}\}\cap S|=1$. We now provide a binary assignment on the variables $x\in A$ the following way: for any $x\in A$, if $w^{x}_{1}\in S$ and $w^{x}_{2}\notin S$, then put $(x,\neg x)=(1,0)$; and if $w^{x}_{1}\notin S$ and $w^{x}_{2}\in S$, then put $(x,\neg x)=(0,1)$. Clearly, this assignment on $\psi$ is a valid one since, for each $x\in A$, exactly one of $x$ and $\neg x$ is assigned $1$ and the other 00. To now prove that this assignment is also a satisfying one on $\psi$, we simply note that, for each clause $y\in B$, in order for the set $S$ to full-separate $u^{y}_{1}$ and $u^{y}_{3}$, there exists a variable $x\in A$ such that either $u^{y}_{1}w^{x}_{1}\in E(G^{\psi})$ and $w^{x}_{1}\in S$ or $u^{y}_{1}w^{x}_{2}\in E(G^{\psi})$ and $w^{x}_{2}\in S$. Therefore, by construction of the graph $G^{\psi}$, either $x$ is a variable in the clause $y$ with $(x,\neg x)=(1,0)$ or $\neg x$ is a variable in the clause $y$ with $(x,\neg x)=(0,1)$. Hence, the binary assignment formulated is a satisfying one on $\psi$. ∎
Lemma 3
The following statements are true.
  1. 1.
    If there exists an FTD-code $C$ of $G^{\psi}$ such that $|C|\leq 7n+2m$, then $\psi$ has a satisfying assignment.
  2. 2.
    If there exists an FD-code $C^{\prime}$ of $G^{\psi}$ such that $|C^{\prime}|\leq 7n+2m-1$, then $\psi$ has a satisfying assignment.
Proof
Let $C$ be an FTD-code of $G^{\psi}$ such that $|C|\leq 7n+2m$. Then, we show that $\psi$ has a satisfying assignment. Therefore, by Lemma A.2, we only need to show that |$|\{w^{x}_{1},w^{x}_{2}\}\cap C|=1$ (note that |$|\{w^{x}_{1},w^{x}_{2}\}\cap C^{\prime}|\geq 1$ in order for $C$ to full-separate the vertices $v^{x}_{1},v^{x}_{3}$). If, on the contrary, |$|\{w^{x^{\star}}_{1},w^{x^{\star}}_{2}\}\cap C|=2$ for some $x^{\star}\in A$, then, by Lemmas A.1a and b, this implies that
$7n+$ $2m\geq|C|$
$=\sum_{x\in A}|V(G^{x})\cap C|+\sum_{y\in B}|V(P^{y})\cap C|$
$=\sum_{x\in A}|(V(G^{x})\setminus\{w^{x}_{1},w^{x}_{2}\})\cap C|+\sum_{y\in B}|V(P^{y})\cap C|+\sum_{x\in A}|\{w^{x}_{1},w^{x}_{2}\}\cap C|$
$\geq 6n+2m+|\{w^{x^{\star}}_{1},w^{x^{\star}}_{2}\}\cap C|+\sum_{x\in A\setminus\{x^{\star}\}}|\{w^{x}_{1},w^{x}_{2}\}\cap C|\geq 7n+2m+1.$
However, this is a comtradiction. Therefore, we must have |$|\{w^{x}_{1},w^{x}_{2}\}\cap C|=1$.
We now let $C^{\prime}$ be an FD-code of $G^{\psi}$ such that $|C^{\prime}|\leq 7n+2m-1$ and show that $\psi$ has a satisfying assignment. Then, again by Lemma A.2, we only need to show that |$|\{w^{x}_{1},w^{x}_{2}\}\cap C^{\prime}|=1$. If, on the contrary, |$|\{w^{x^{\star}}_{1},w^{x^{\star}}_{2}\}\cap C^{\prime}|=2$ for some $x^{\star}\in A$, then, by Lemma A.1c, we have $|V(G^{x})\setminus\{w^{x}_{1},w^{x}_{2}\}\cap C^{\prime}|\geq 6$ for all $x\in A$. Again, since |$|\{w^{x}_{1},w^{x}_{2}\}\cap C^{\prime}|\geq 1$ in order for $C$ to full-separate the vertices $v^{x}_{1},v^{x}_{3}$, we must have $|V(G^{x})\cap C^{\prime}|\geq 7$ for all $x\in A$. This implies that
$7n+2m-1\geq|C^{\prime}|$ $=\sum_{x\in A}|V(G^{x})\cap C^{\prime}|+\sum_{y\in B}|V(P^{y})\cap C^{\prime}|\geq 7n+2m.$
This is again a contradiction. Therefore, we must have |$|\{w^{x}_{1},w^{x}_{2}\}\cap C^{\prime}|=1$. ∎
4 Proofs of results in Section 4: FD- and FTD-numbers of some graph families
Paths and cycles
Lemma A.3
Let $G$ be is either a path $P_{n}$ for $n\geq 4$ or a cycle $C_{n}$ for $n\geq 5$. Then, we have $\gamma^{\rm{FD}}(G)=\gamma^{\rm{FTD}}(G)$.
Proof
The graph $G$ has maximum degree $2$. Let $G=(V(G),E(G))=(V,E)$. On the contrary, let us assume that $\gamma^{\rm{FD}}(G)\neq\gamma^{\rm{FTD}}(G)$. Moreover, let $C$ be a minimum FD-code of $G$. Therefore, we have $|C|=\gamma^{\rm{FD}}(G)$. By Observation A.2e, there exists exactly one vertex $w$ of $G$ such that $N(w)\cap C=\emptyset$. If $C$ does not full-separate any pair of distinct vertices of $G$ only by $w$, then, for any neighbor $w^{\prime}$ of $w$ in $G$, the set $C^{\prime}=(C\setminus\{w\})\cup\{w^{\prime}\}$ is an FTD-code of $G$ such that $|C^{\prime}|=|C|=\gamma^{\rm{FD}}(G)$. This implies that $\gamma^{\rm{FTD}}(G)\leq\gamma^{\rm{FD}}(G)$ and hence, $\gamma^{\rm{FD}}(G)=\gamma^{\rm{FTD}}(G)$, contrary to our assumption. Therefore, the set $C$ must full-separate some distinct $u,v\in V$ uniquely by $w$. This implies, without loss of generality, that $u\in N(w)$ and $v\notin N[w]$. Moreover, since $N(w)\cap C=\emptyset$ and $v\neq w$, by Observation A.2(c), we must have $N(v)\cap C\neq\emptyset$. Therefore, let $x\in N(v)\cap C$. Then, $x\notin N[w]$. Now, since $C$ full-separates $u$ and $v$ uniquely by $w$, we must have $x\in N(u)$ as well. Since $u\notin C$, in order for $C$ to full-separate $v$ and $x$, there must exist another vertex $y\in(N(v)\setminus N(x))\cap C$ (notice that $y$ cannot be a neighbor of $x$ since $\{u,v\}\subset N(x)$ and $deg(x)\leq 2$). Moreover, $uy\notin E$ since $\{w,x\}\subset N(u)$ and $deg(u)\leq 2$. Thus, the set $C$ full-separates $u$ and $v$ by $y$, a contradiction to the uniqueness of $w$. Hence, we must have $\gamma^{\rm{FD}}(G)=\gamma^{\rm{FTD}}(G)$. ∎
Observation A.3
Let $G=(V,E)$ be either a path $P_{n}$ for $n\geq 4$ or a cycle $C_{n}$ for $n\geq 5$. Let $C$ be an FTD-code of $G$. Then, for any vertex $v\in V$, there exists a vertex subset $S_{v}\subset C$ of $G$ such that $v\in S_{v}$, $|S_{v}|\geq 4$ and $G[S_{v}]$ is the path $P_{|S_{v}|}$.
Proof
Let $V=\{v_{1},v_{2},\ldots,v_{n}\}$ such that $v_{i}v_{i+1}\in E$ for all $i\in[n-1]$ and $v_{n}vi\in E$ if $G\cong C_{n}$. Let $v=v_{i}\in C$ for some $i\in[n]$. In order for $C$ to total-dominate $v_{i}$, we must have $v_{i+1}\in C$, say, without loss of generality. Next, in order for $C$ to full-separate $v_{i}$ and $v_{i+1}$, we must have $v_{i+2}\in C$, say, without loss of generality. Again, in order for $C$ to full-separate $v_{i}$ and $v_{i+2}$, we must have $v_{i+3}\in C$, say, without loss of generality. Thus, the result follows by taking $S=\{v_{i},v_{i+1},v_{i+2},v_{i+3}\}$. ∎
Lemma A.4
Let $G$ be either a path $P_{n}$ for $n\geq 4$ or a cycle $C_{n}$ for $n\geq 5$. Let $n=6q+r$, where $q$ and $r$ are non-negative integers with $r\in[0,5]$. Then we have
$\gamma^{\rm{FTD}}(G)\leq\begin{cases}4q+r,&\text{if $r\in[0,4]$;}\\ 4q+4,&\text{if $r=5$}.\end{cases}$
Proof
Let $V=V(G)=\{v_{1},v_{2},\ldots,v_{n}\}$ and that $n=6q+r$, where $r\in[0,5]$. If $q=0$, the graph $G$ is either a $4$-path or a $5$-path or a $5$-cycle. In all the three cases, it can be checked that the set $\{v_{1},v_{2},v_{3},v_{4}\}$ is an FTD-code of $G$. Thus, in these cases, the result holds. For the rest of this proof, therefore, we assume that $n\geq 6$, that is, $q\geq 1$. We now construct a vertex subset $C$ of $G$ by including in $C$ the vertices $v_{6k-4},v_{6k-3},v_{6k-2},v_{6k-1}$ for all $k\in[q]$, the vertices $v_{6q},v_{6q+1},\ldots,v_{6q+r-1}$ if $r\in[4]$ and the vertices $v_{6q+1},v_{6q+2},v_{6q+3},v_{6q+4}$ if $r=5$ (see Figure 5 for an example with the code vertices marked in black). Notice that, by construction, the vertex subset $C$ is a total-dominating set of $G$. Therefore, by Observation A.2f to show that $C$ is an FTD-code of $G$, it is enough to show that $C$ full-separates every pair of distinct $u,v\in V$ such that $d(u,v)\leq 2$. Now, one can observe that for every pair $v_{i},v_{j}$ of distinct vertices in $G$ such that $i<j\leq i+2$ (that is $d_{G}(v_{i},v_{j})\leq 2$), either $v_{i-1}$ or $v_{j+1}$ (or both) is in $C$. Thus, all vertex pairs $v_{i},v_{j}\in V$ with $d_{G}(v_{i},v_{j})\leq 2$ are full-separated by $C$. This implies that $C$ is an FTD-code of $G$. Hence, by counting the order of $C$ from its contsruction, the result follows by $\gamma^{\rm{FTD}}(G)\leq|C|$. ∎
Theorem 4.6
Let $G$ be either a path $P_{n}$ for $n\geq 4$ or a cycle $C_{n}$ for $n\geq 5$. Moreover, let $n=6q+r$ for non-negative integers $q$ and $r\in[0,5]$. Then, we have
$\gamma^{\rm{FD}}(G)=\gamma^{\rm{FTD}}(G)=\begin{cases}4q+r,&\text{if $r\in[0,4]$};\\ 4q+4,&\text{if $r=5$}.\end{cases}$
Proof
Let $V=V(G)=\{v_{1},v_{2},\ldots,v_{n}\}$ such that $v_{i}v_{i+1}\in E$ for all $i\in[n-1]$ and $v_{n}vi\in E$ if $G\cong C_{n}$.. By Lemma A.3, it is enough to prove the result only for $\gamma^{\rm{FTD}}(G)$. To begin with, let us assume that $G$ is a path $P_{n}$ for $n\geq 4$. Then, by Observation A.2d, an FTD-code of $G$ is also an OTD-code of $G$. Therefore, we have
$\gamma^{\rm{FTD}}(P_{n})\geq\gamma^{\rm{OTD}}(P_{n})=\begin{cases}4q+r,&\text{if $r\in[0,4]$}\\ 4q+4,&\text{if $r=5$}.\end{cases}$
The last inequality is due to Seo and Slater [21]. Therefore, by Lemma A.4, the result holds for the FTD-numbers of paths on $n\geq 4$ vertices.
We now prove the result for $\gamma^{\rm{FTD}}(G)$ when $G$ is a cycle $C_{n}$ for $n\geq 5$. Bianchi et al. had shown in [6] that
$\gamma^{\rm{OTD}}(C_{n})=\begin{cases}4q+r,&\text{if $r=0,1,2,4$}\\ 4q+2,&\text{if $r=3$}\\ 4q+4,&\text{if $r=5$}.\end{cases}$ (1)
Then again, by Lemma A.4, the result for $\gamma^{\rm{FTD}}(C_{n})$ holds for all $n=6q+r$ except for $r=3$. Therefore, we next analyse only the $(6q+3)$-cycles. So, let $n=6q+3$ for any positive integer $q$. Using Lemma A.4, we notice that it is enough to prove that $\gamma^{\rm{FTD}}(C_{n})\geq 4q+3$. So, let us assume on the contrary that there exists an FTD-code $C$ of $G$ such that $|C|\leq 4q+2$.
Claim
For some $w\in V$, there exists a vertex subset $S_{w}\subset C$ of $G$ (as in Observation A.3) such that $w\in S_{u}$, $|S_{w}|\geq 5$ and $G[S_{w}]$ is the path $P_{|S_{w}|}$.
Proof (of Claim)
Without loss of generality, let us assume that, for every $v\in V$, the sets $S_{v}$ as in Observation A.3 are maximal. Contray to the claim, let us assume that $|S_{v}|\leq 4$ for all $v\in V$. Then, again using Observation A.3, we have $|S_{v}|=4$, for all $v\in V$. Define $u\sim v$ if $S_{u}=S_{v}$. For some $i<j$ and $v_{i}\not\sim v_{j}$, if we have $\max\{i^{\prime}:v_{i^{\prime}}\in S_{v_{i}}\}=\min\{j^{\prime}:v_{j^{\prime}}\in S_{v_{j}}\}$, then it implies that $|S_{v_{i}}|\geq 8$, a contradiction. Hence, for every $i<j$ and $v_{i}\not\sim v_{j}$, there must be a vertex $v_{k}\in V\setminus C$ such that $\max\{i^{\prime}:v_{i^{\prime}}\in S_{v_{i}}\}<k<\min\{j^{\prime}:v_{j^{\prime}}\in S_{v_{j}}\}$. This implies that $|C|=4q^{\prime}$, where $q^{\prime}\in[q]$. Now, for any $i<j$ and $v_{i}\not\sim v_{j}$, let $v_{k},v_{k+1},\ldots,v_{k+k^{\prime}}\in V\setminus C$ such that $\max\{i^{\prime}:v_{i^{\prime}}\in S_{v_{i}}\}<k<k+1<\ldots<k+k^{\prime}<\min\{j^{\prime}:v_{j^{\prime}}\in S_{v_{j}}\}$. Then we must have $k^{\prime}\in\{0,1\}$, or else, the vertex $v_{k+1}$ is not total-dominated by $C$, a contradiction. Therefore, by counting, we must have $n\leq 6q^{\prime}\leq 6q$ which is a contradiction since $n=6q+3$. This proves our claim.
Now, let $S=S_{w}$ and $v_{i},v_{i+1}\in S$. Then, let $G^{\prime}$ be the $(6q+2)$-cycle obtained by contracting the edge $v_{i}v_{i+1}$. Let $V^{\prime}=V(G^{\prime})=\{u_{1},u_{2},\ldots,u_{6q+2}\}$, where $u_{j}=v_{j}$ for $j\in[i]$ and $u_{j}=v_{j+1}$ for $j\in[i+1,6q+2]$. Moreover, let $C^{\prime}=\{u_{j}:v_{j}\in C\}$. . Then $C^{\prime}$ is also a total-dominating set of $G^{\prime}$. Moreover, we have $|S_{u_{j}}|\geq 4$ for all $j\in[6q+2]$. For any two vertices $u_{j},u_{j^{\prime}}\in V^{\prime}$, where $j<j^{\prime}$ and $j^{\prime}-j\leq 2$, we can observe that either $u_{j-1}$ or $u_{j^{\prime}+1}$ (or both) is in $C^{\prime}$. Hence, by Observation A.2f, the set $C^{\prime}$ is an FTD-code of $G^{\prime}$. However, this is a contradiction to Equation (1), since now we have $\gamma^{\rm{FTD}}(C_{6q+2})\leq|C^{\prime}|=4q+1$. Hence, $|C|\geq 4q+3$ and this proves the result. ∎
In order to calculate the X-numbers of the other studied graphs $G$, we are going to construct the X-hypergraphs $\mathcal{H}_{\textsc{X}}(G)$ and to determine their covering numbers $\tau(\mathcal{H}_{\textsc{X}}(G))$. In this context, it was observed in, for example, [1, 2, 3, 8] that $\mathcal{H}_{\textsc{X}}(G)=(V,\mathcal{F}_{\textsc{X}})$ may contain redundant hyperedges. In fact, if there are two hyperedges $F,F^{\prime}\in\mathcal{F}_{\textsc{X}}$ with $F\subseteq F^{\prime}$, then $F\cap C\neq\emptyset$ also implies $F^{\prime}\cap C\neq\emptyset$ for every $C\subseteq V$. Thus, $F^{\prime}$ is redundant as $(V,\mathcal{F}_{\textsc{X}}-\{F^{\prime}\})$ suffices to encode the X-codes of $G$. Hence, only non-redundant hyperedges of $\mathcal{H}_{\textsc{X}}(G)$ need to be considered in order to determine $\tau(\mathcal{H}_{\textsc{X}}(G))$ and thus $\gamma^{\rm{X}}(G)$.
Half-graphs.
Theorem 4.7
For a half-graph $B_{k}$ with $k\geq 3$, we have $\gamma^{\rm{FD}}(B_{k})=2k-1$.
Proof
Consider a half-graph $B_{k}=(U\cup W,E)$ with vertices in $U=\{u_{1},\ldots,u_{k}\}$, $W=\{w_{1},\ldots,w_{k}\}$ and edges $u_{i}w_{j}$ if and only if $i\leq j$ and $k\geq 3$.
In order to calculate $\gamma^{\rm{FD}}(B_{k})$, we first construct $\mathcal{H}_{\textsc{FD}}(B_{k})$. From the definition of FD-codes and Theorem 2.1, we know that $\mathcal{H}_{\textsc{FD}}(B_{k})$ is composed of the closed neighborhoods
  • $N[u_{i}]=\{u_{i}\}\cup\{w_{i},\ldots,w_{k}\}$ for all $u_{i}\in U$,
  • $N[w_{j}]=\{u_{1},\ldots,u_{j}\}\cup\{w_{j}\}$ for all $w_{j}\in W$,
the symmetric differences of closed neighborhoods of adjacent vertices
  • $N[u_{i}]\bigtriangleup N[w_{j}]=\{u_{1},\ldots,u_{i-1},u_{i+1}\ldots,u_{j}\}\cup\{w_{i},\ldots,w_{j-1},w_{j+1}\ldots,w_{k}\}$ for $i<j$,
  • $N[u_{i}]\bigtriangleup N[w_{i}]=\{u_{1},\ldots,u_{i-1}\}\cup\{w_{i+1}\ldots,w_{k}\}$
and the symmetric differences of open neighborhoods of non-adjacent vertices
  • $N(u_{i})\bigtriangleup N(u_{j})=\{w_{i},\ldots,w_{j-1}\}$ for $i<j$,
  • $N(w_{i})\bigtriangleup N(w_{j})=\{u_{i+1}\ldots,u_{j}\}$ for $i<j$,
  • $N(u_{i})\bigtriangleup N(w_{j})=\{w_{i},\ldots,w_{k}\}\cup\{u_{1},\ldots,u_{j}\}$ for $i>j$.
In particular, we see that
  • $N(u_{i})\bigtriangleup N(u_{i+1})=\{w_{i}\}$ for $1\leq i<k$,
  • $N(w_{i})\bigtriangleup N(w_{i+1})=\{u_{i+1}\}$ for $1\leq i<k$, and
  • $N(u_{k})\bigtriangleup N(w_{1})=\{w_{k}\}\cup\{u_{1}\}$
holds. Therefore, for $k\geq 3$, all neighborhoods and all other symmetric differences are redundant hyperedges (note that this is not the case for $k=3$ as then $N[u_{1}]\bigtriangleup N[w_{1}]=\{w_{2}\}$ and $N[u_{2}]\bigtriangleup N[w_{2}]=\{u_{1}\}$ holds and makes $N(u_{2})\bigtriangleup N(w_{1})=\{w_{2}\}\cup\{u_{1}\}$ redundant). This clearly shows that $U\cup W\setminus\{w_{k}\}$ and $U\setminus\{u_{1}\}\cup W$ are the only two minimum FD-codes of the half-graph $B_{k}$ with $k\geq 3$, hence $\gamma^{\rm{FD}}(B_{k})=2k-1$ indeed follows. ∎
Headless spiders.
Theorem 4.8
For a thin headless spider $H_{k}=(Q\cup S,E)$ with $k\geq 4$, we have $\gamma^{\rm{FD}}(H_{k})=2k-2$ and $\gamma^{\rm{FTD}}(H_{k})=2k-1$.
Proof
[…]
  • $N[q_{i}]=Q\cup\{s_{i}\}$ or $N(q_{i})=Q\setminus\{q_{i}\}\cup\{s_{i}\}$ for all $q_{i}\in Q$,
  • $N[s_{i}]=\{q_{i},s_{i}\}$ or $N(s_{i})=\{q_{i}\}$ for all $s_{i}\in S$,
the symmetric differences of closed neighborhoods of adjacent vertices, that is,
  • $N[s_{i}]\bigtriangleup N[q_{i}]=Q\setminus\{q_{i}\}$ for $1\leq i\leq k$,
  • $N[q_{i}]\bigtriangleup N[q_{j}]=\{s_{i},s_{j}\}$ for $1\leq i<j\leq k$,
and the symmetric differences of closed or open neighborhoods of non-adjacent vertices, that is,
  • $N[s_{i}]\bigtriangleup N[q_{j}]=Q\setminus\{q_{i}\}\cup\{s_{i},s_{j}\}$ or $N(s_{i})\bigtriangleup N(q_{j})=Q\setminus\{q_{i},q_{j}\}\cup\{s_{j}\}$ for $i\neq j$,
  • $N[s_{i}]\bigtriangleup N[s_{j}]=\{q_{i},q_{j}\}\cup\{s_{i},s_{j}\}$ or $N(s_{i})\bigtriangleup N(s_{j})=\{q_{i},q_{j}\}$ for $1\leq i<j\leq k$.
$\mathcal{H}_{\textsc{FD}}(H_{k})$ is composed of the closed neighborhoods and the symmetric differences of closed neighborhoods of adjacent vertices as well as of open neighborhoods of non-adjacent vertices. In particular, we see that
  • $N[s_{i}]=\{q_{i},s_{i}\}$ for $1\leq i\leq k$,
  • $N[q_{i}]\bigtriangleup N[q_{j}]=\{s_{i},s_{j}\}$ for $1\leq i<j\leq k$,
  • $N(s_{i})\bigtriangleup N(s_{j})=\{q_{i},q_{j}\}$ for $1\leq i<j\leq k$
belong to $\mathcal{H}_{\textsc{FD}}(H_{k})$. Therefore, $N[q_{i}]$ for all $q_{i}\in Q$, $N[s_{i}]\bigtriangleup N[q_{i}]$ for $1\leq i\leq k$, and $N(s_{i})\bigtriangleup N(q_{j})$ for $i\neq j$ are redundant for $k\geq 4$ (note that $N(s_{i})\bigtriangleup N(q_{j})$ is not redundant if $k=3$). This shows that $Q\setminus\{q_{i}\}\cup S\setminus\{s_{j}\}$ for $i\neq j$ are the minimum FD-codes of $H_{k}$, hence $\gamma^{\rm{FD}}(H_{k})=2k-2$ follows.
$\mathcal{H}_{\textsc{FTD}}(H_{k})$ is composed of the open neighborhoods and the symmetric differences of closed neighborhoods of adjacent vertices as well as of open neighborhoods of non-adjacent vertices. In particular, we have that
  • $N(s_{i})=\{q_{i}\}$ for $1\leq i\leq k$,
  • $N[q_{i}]\bigtriangleup N[q_{j}]=\{s_{i},s_{j}\}$ for $1\leq i<j\leq k$
belong to $\mathcal{H}_{\textsc{FTD}}(H_{k})$. Therefore, $N(q_{i})$ for all $q_{i}\in Q$, $N[s_{i}]\bigtriangleup N[q_{i}]$ for $1\leq i\leq k$, $N(s_{i})\bigtriangleup N(q_{j})$ for $i\neq j$ and $N(s_{i})\bigtriangleup N(s_{j})$ for $1\leq i<j\leq k$ are redundant (even for $k\geq 3$). We conclude that $Q\cup S\setminus\{s_{i}\}$ for $1\leq i\leq k$ are the minimum FTD-codes of $H_{k}$ and $\gamma^{\rm{FTD}}(H_{k})=2k-1$ holds accordingly. ∎
Theorem 4.9
For a thick headless spider $\overline{H}_{k}=(Q\cup S,E)$ with $k\geq 4$, we have $\gamma^{\rm{ITD}}(\overline{H}_{k})=k+1$ and $\gamma^{\rm{FD}}(\overline{H}_{k})=\gamma^{\rm{FTD}}(\overline{H}_{k})=2k-2$.
Proof
[…]
  • $N[s_{i}]=Q\setminus\{q_{i}\}\cup\{s_{i}\}$ or $N(s_{i})=Q\setminus\{q_{i}\}$ for all $s_{i}\in S$,
  • $N[q_{i}]=Q\cup S\setminus\{s_{i}\}$ or $N(q_{i})=Q\setminus\{q_{i}\}\cup S\setminus\{s_{i}\}$ for all $q_{i}\in Q$,
the symmetric differences of closed neighborhoods of adjacent vertices, that is,
  • $N[s_{i}]\bigtriangleup N[q_{j}]=\{q_{i}\}\cup S\setminus\{s_{i},s_{j}\}$ for $i\neq j$,
  • $N[q_{i}]\bigtriangleup N[q_{j}]=\{s_{i},s_{j}\}$ for $i\neq j$,
and the symmetric differences of closed or open neighborhoods of non-adjacent vertices, that is,
  • $N[s_{i}]\bigtriangleup N[q_{i}]=\{q_{i}\}\cup S$ or $N(s_{i})\bigtriangleup N(q_{i})=S\setminus\{s_{i}\}$ for $1\leq i\leq k$,
  • $N[s_{i}]\bigtriangleup N[s_{j}]=\{q_{i},q_{j}\}\cup\{s_{i},s_{j}\}$ or $N(s_{i})\bigtriangleup N(s_{j})=\{q_{i},q_{j}\}$ for $i\neq j$.
$\mathcal{H}_{\textsc{FD}}(\overline{H}_{k})$ is composed of the closed neighborhoods and the symmetric differences of closed neighborhoods of adjacent vertices as well as of open neighborhoods of non-adjacent vertices. In particular, we see that
  • $N[q_{i}]\bigtriangleup N[q_{j}]=\{s_{i},s_{j}\}$ for $1\leq i<j\leq k$,
  • $N(s_{i})\bigtriangleup N(s_{j})=\{q_{i},q_{j}\}$ for $1\leq i<j\leq k$
belong to $\mathcal{H}_{\textsc{FD}}(\overline{H}_{k})$. Therefore, all neighborhoods and the symmetric differences $N[s_{i}]\bigtriangleup N[q_{j}]$ for $i\neq j$ and $N(s_{i})\bigtriangleup N(q_{i})$ for $1\leq i\leq k$ are redundant for $k\geq 4$ (note that $N[s_{i}]\bigtriangleup N[q_{j}]$ are not redundant if $k=3$). This shows that $Q\setminus\{q_{i}\}\cup S\setminus\{s_{j}\}$ for $1\leq i,j\leq k$ are the minimum FD-codes of $\overline{H}_{k}$, hence $\gamma^{\rm{FD}}(\overline{H}_{k})=2k-2$ follows.
$\mathcal{H}_{\textsc{FTD}}(\overline{H}_{k})$ is composed of the open neighborhoods and the symmetric differences of closed neighborhoods of adjacent vertices as well as of open neighborhoods of non-adjacent vertices. Again, we have that
  • $N[q_{i}]\bigtriangleup N[q_{j}]=\{s_{i},s_{j}\}$ for $1\leq i<j\leq k$,
  • $N(s_{i})\bigtriangleup N(s_{j})=\{q_{i},q_{j}\}$ for $1\leq i<j\leq k$
belong to $\mathcal{H}_{\textsc{FTD}}(\overline{H}_{k})$. Therefore, all neighborhoods and the other symmetric differences are ag redundanaint for $k\geq 4$. We conclude that $Q\setminus\{q_{i}\}\cup S\setminus\{s_{j}\}$ for $1\leq i,j\leq k$ are also the minimum FTD-codes of $\overline{H}_{k}$ and thus also $\gamma^{\rm{FTD}}(\overline{H}_{k})=2k-2$ holds.
$\mathcal{H}_{\textsc{ITD}}(\overline{H}_{k})$ is composed of the open neighborhoods and the symmetric differences of closed neighborhoods of distinct vertices. In particular, we see that
  • $N(s_{i})=Q\setminus\{q_{i}\}$ for $1\leq i\leq k$,
  • $N[q_{i}]\bigtriangleup N[q_{j}]=\{s_{i},s_{j}\}$ for $1\leq i<j\leq k$
belong to $\mathcal{H}_{\textsc{ITD}}(\overline{H}_{k})$. Thus, $N(q_{i})$ for all $q_{i}\in Q$ and all other symmetric differences are redundant for $k\geq 4$ (note that $N[s_{i}]\bigtriangleup N[q_{j}]$ are not redundant if $k=3$). This implies that two vertices from $Q$ and all but one vertex from $S$ have to be taken for every minimum ITD-code of $\overline{H}_{k}$, therefore $\gamma^{\rm{FTD}}(\overline{H}_{k})=k+1$ follows. ∎